Determine the total energy of an electron traveling at 0.98c.


The answer is 2.6 MeV


I need to understand the steps.
Answer #1

it is a relativistic case.
Eo = mo*c^2 = 0.5 Mev


m = mo/sqrt(1- (v/c)^2)

E = m*c^2 = 2.55 MeV
Answer #2
First the Mass of electron is 9.11*10^-31 E=MC² E=9.11*10^-31(3*10^8)= .....J Convert J to ev Convert ev to Mev Then use the law M°=m/sqrt(1-(v/c)^2) Total energy is 2.51Mev
source: I am chemist trus me bro enjoy
answered by: anonymous